OpenStudy (anonymous):

In a shop there are 20 customers, 18 of whom make a purchase. If three customers are selected, one at a time, at random, what is the probability that all will make a purchase?

OpenStudy (anonymous):

\[\frac{18}{20}\times \frac{17}{19}\times \frac{16}{18}\]